# Do I Know This Entity? Knowledge Awareness and Hallucinations in Language Models

### v1

@@ Abstract
Hallucinations in large language models are a widespread problem, yet the mechanisms behind whether models will hallucinate are poorly understood, limiting our ability to solve this problem. Using sparse autoencoders as an interpretability tool, we discover that a key part of these mechanisms is entity recognition, where the model detects if an entity is one it can recall facts about. Sparse autoencoders uncover meaningful directions in the representation space, these detect whether the model recognizes an entity, e.g. detecting it doesn’t know about an athlete or a movie. This suggests that models can have self-knowledge: internal representations about their own capabilities. These directions are causally relevant: capable of steering the model to refuse to answer questions about known entities, or to hallucinate attributes of unknown entities when it would otherwise refuse. We demonstrate that despite the sparse autoencoders being trained on the base model, these directions have a causal effect on the chat model’s refusal behavior, suggesting that chat finetuning has repurposed this existing mechanism. Furthermore, we provide an initial exploration into the mechanistic role of these directions in the model, finding that they disrupt the attention of downstream heads that typically move entity attributes to the final token.
@@ 1 Introduction
Large Language Models (LLMs) have remarkable capabilities (Radford et al., 2019; Brown et al., 2020; Hoffmann et al., 2022; Chowdhery et al., 2023) yet have a propensity to hallucinate: generating text that is fluent but factually incorrect or unsupported by available information (Ji et al., 2023; Minaee et al., 2024). This significantly limits their application in real-world settings where factuality is crucial, such as healthcare. Despite the prevalence and importance of this issue, the mechanistic understanding of whether LLMs will hallucinate on a given prompt remains limited. While there has been much work interpreting factual recall (Geva et al., 2023; Nanda et al., 2023; Chughtai et al., 2024), it has mainly focused on the mechanism behind recalling known facts, not on hallucinations or refusals to answer, leaving a significant gap in our understanding.
Language models can produce hallucinations due to various factors, including flawed data sources or outdated factual knowledge (Huang et al., 2023). However, an important subset of hallucinations occurs when models are prompted to generate information they don’t possess. We operationalize this phenomenon by considering queries about entities of different types (movies, cities, players, and songs). Given a question about an unknown entity, the model either hallucinates or refuses to answer. In this work, we find linear directions in the representation space that encode a form of self-knowledge: assessing their own knowledge or lack thereof regarding specific entities. These directions are causally relevant for whether it refuses to answer. We note that the existence of this kind of self-knowledge does not necessarily imply the existence of other forms of self-knowledge, and may be specific to the factual recall mechanism.
We find these directions using Sparse Autoencoders (SAEs) (Bricken et al., 2023; Cunningham et al., 2023). SAEs are an interpretability tool for finding a sparse, interpretable decomposition of model representations. They are motivated by the Linear Representation Hypothesis (Park et al., 2023; Mikolov et al., 2013): that interpretable properties of the input (features) such as sentiment (Tigges et al., 2023) or truthfulness (Li et al., 2023; Zou et al., 2023) are encoded as linear directions in the representation space, and that model representations are sparse linear combinations of these directions. We use Gemma Scope (Lieberum et al., 2024), which offers a suite of SAEs trained on every layer of Gemma 2 models (Team et al., 2024), and find internal representations of self-knowledge in Gemma 2 2B and 9B.
@@
Using sparse autoencoders (SAEs) we discover directions in the representation space on the final token of an entity, detecting whether the model can recall facts about the entity, a form of self-knowledge about its own capabilities.
@@ 2 Sparse Autoencoders
Dictionary learning (Olshausen & Field, 1997) offers a powerful approach for disentangling features in superposition. Sparse Autoencoders (SAEs) have proven to be effective for this task (Sharkey et al., 2022; Bricken et al., 2023). SAEs project model representations into a larger dimensional space . In this work, we use the SAEs from Gemma Scope (Lieberum et al., 2024)111We use the default sparsity for each layer, the ones available in Neuronpedia (Lin & Bloom, 2024)., which use the JumpReLU SAE architecture (Rajamanoharan et al., 2024), which defines the function
@@
Equation 1 shows that the model representation can be approximately reconstructed by a linear combination of the SAE decoder latents, which often represent monosemantic features (Cunningham et al., 2023; Bricken et al., 2023; Templeton et al., 2024; Gao et al., 2024). By incorporating a sparsity penalty into the training loss function, we can constrain this reconstruction to be a sparse linear combination, thereby enhancing interpretability:
@@ 3 Methodology
We then categorize entities into ‘known’ or ‘unknown’. Known entities are those where the model gets at least two attributes correct, while unknown are where it gets them all wrong, we discard any in-between. To measure correctness we use fuzzy string matching222https://github.com/seatgeek/thefuzz.. See Appendix A for a description of the process. Finally, we split the entities into train/validation/test (50%, 10%, 40%) sets.
@@ 4 Sparse Autoencoders Uncover Entity Recognition Directions
We also examine the level of generality of the latents by measuring their minimum separation score across entity types (): players, song, cities and movies. A high minimum separation score indicates that a latent performs robustly across entity types, suggesting strong generalization capabilities. For this purpose, for each layer () we compute , and similarly for unknown entities. The increasing trend shown in the MaxMin (red) line in Figure 2 for Gemma 2 2B and in Figure 9, Appendix D for Gemma 2 9B suggests that more generalized latents—those that distinguish between known and unknown entities across various entity types—are concentrated in these intermediate layers. This finding points to a hierarchical organization of entity representation within the model, with more specialized, worse quality, latents in earlier layers and more generalized, higher quality entity-type-agnostic features emerging in the middle layers.
@@
Table 1 demonstrates the activation patterns of the Gemma 2 2B topmost known entity latent on prompts with well-known entities (left), and the patterns for the topmost unknown entity latent (right), firing across entities of different types that cannot be recognized. In Appendix B we provide the activations of these latents on sentences containing a diverse set of entity types, suggesting that indeed they are highly general. In the following sections, we explore how these latents influence the model’s overall behavior.
@@ 5 Entity Recognition Directions Causally Affect Knowledge Refusal
and we define a set of common knowledge refusal completions and detect if any of these occur with string matching, e.g. ‘Unfortunately, I don’t have access to real-time information…’. Gemma 2 includes both a base model, and a fine-tuned chat (i.e. instruction tuned) model. In Section 4 we found the entity recognition latents by studying the base model, but here focus on the chat model, as they have been explicitly fine-tuned to perform knowledge refusal where appropriate (Team et al., 2024)333The Gemma 2 technical report (Team et al., 2024) mentions “including subsets of data that encourage refusals to minimize hallucinations improves performance on factuality metrics”. This pattern is consistent with recent language models, such as Llama 3.1 (Dubey et al., 2024), where the explicit finetuning process for knowledge refusal has been documented., and the factuality of chat models is highly desirable.
@@
We hypothesize that entity recognition directions could be used by chat models to induce knowledge refusal. To evaluate this, we use a test set sample of 100 questions about unknown entities, and measure the number of times the model refuses by steering (as in Equation 4) with the entity recognition latents the last token of the entity and the following end-of-instruction-tokens.444We use a validation set to select an appropriate steering coefficient . In Appendix G we show generations of Gemma 2B IT with different steering coefficients. We select , which corresponds to around two times the norm of the residual stream in the layers where the entity recognition latents are present (Appendix E). Figure 3 (left) illustrates the original model refusal rate (blue bar), showing some refusal across entity types. We see that the entity recognition SAE latents found in the base model transfer to the chat model and, by increasing the known entity latent activation we are able to reduce refusal in three entity types. Conversely, increasing the unknown entity latent induces almost 100% refusal across all entity types. We also include an Orthogonalized model baseline, which consists of doing weight orthogonalization (Arditi et al., 2024) on every matrix writing to the residual stream. Weight orthogonalization modifies each row of a weight matrix to make it perpendicular to a specified direction vector . This is achieved by subtracting the component of each row that is parallel to :
@@
By doing this operation on every output matrix in the model we ensure no component is able to write into that direction. The resulting orthogonalized model exhibits a large reduction in refusal responses, suggesting this direction plays a crucial role in the model’s knowledge refusal behavior. We also include the average refusal rate after steering with 10 differents random latents, using the same configuration (layer and steering coefficient) that the known and unknown entity latents respectively. Additional analysis of the Gemma 2 9B model, detailed in Section F, reveals similar patterns, albeit with less pronounced effects compared to the 2B model.
@@ 6 Mechanistic Analysis
In the previous section, we saw that entity recognition latents had a causal effect on knowledge refusal. Here, we look at how they affect the factual recall mechanism (aka circuit) in prompts of the format of Example 5. This has been well studied before on other language models (Nanda et al., 2023; Geva et al., 2023; Meng et al., 2022a). We replicate the approach of Nanda et al. (2023) on Gemma 2 2B and 9B and find a similar circuit. Namely, early attention heads merge the entity’s name into the last token of the entity, and downstream attention heads extract relevant attributes from the entity and move them to the final token position (Figure 4 (a, b)), this pattern holds across various entity types and model sizes (Appendix I and Appendix J). To do the analysis, we perform activation patching (Geiger et al., 2020; Vig et al., 2020; Meng et al., 2022a) on the residual streams and attention heads’ outputs (see Appendix H for a detailed explanation on the method). We use the denoising setup (Heimersheim & Nanda, 2024), where we patch representations from a clean run (with a known entity) and apply it over the run with a corrupted input (with an unknown entity).555We show the proportion of logit difference recovered after each patch in Figure 4 (a). A recovered logit difference of 1 indicates that the prediction after patching is the same as the original prediction in the clean run.
@@
Expanding on the findings of Yuksekgonul et al. (2024), who established a link between prediction accuracy and attention to the entity tokens, our study reveals a large disparity in attention between known and unknown entities, for instance the attribute extraction heads L18H5 and L20H3 (Figure 4 (c)), which are overall relevant across entity types in Gemma 2 2B (see example of attributes extracted by these heads in Appendix L). Notably, attention scores are higher when faced with a known entity. We also observe a causal relationship between the entity recognition latents and the behavior of these attention heads. Steering with the top unknown entity latent reduces the attention to the last token of the entity, even in prompts with a known entity (Figure 4 (d)), while steering with the known entity latent increases the attention scores (Figure 4 (e)). We show in Figure 4 (f) the results of steering with a random vector baseline for comparison, and in Appendix K the results of steering with a random SAE latent. In Appendix M we illustrate the average attention score change to the entity tokens after steering on the residual streams of the last token of the entities in Gemma 2 2B and 9B with the top 3 known and unknown entity latents. The results reveal an increase/decrease attention score across upper layer heads, with the 9B model showing more subtle effects when steered using unknown latents.
@@
7 Uncertainty Directions
Our study focuses on the study of the residual streams before the answer. We hypothesize that end-of-instruction tokens, which always succeed the instruction, may aggregate information about the whole question (Marks & Tegmark, 2023).666This concept was termed by Tigges et al. (2023) as the ‘summarization motif’. We select the token model and use examples such as:
For each entity type and layer with available SAE we extract the representations of the model residual stream, for both correct and mistaken answers, and gather the SAE latent activations. We are interested in seeing whether there are SAE latents that convey information about how unsure or uncertain the model is to answer to a question, but still fails to refuse, giving rise to hallucinations.
To capture subtle variations in model uncertainty, which may be represented even when attributes are correctly recalled, we focus on quantifying differences in activation levels between correct and incorrect responses. For each latent, we compute the t-statistic using two activation samples: for correct responses and for incorrect ones. The t-statistic measures how different the two sample means are from each other, taking into account the variability within the samples:
We use a pre-trained SAE for the 13th layer (out of 18) of Gemma 2B IT777https://huggingface.co/jbloom/Gemma-2b-IT-Residual-Stream-SAEs. We note that Gemma Scope doesn’t provide SAEs for Gemma 2 2B IT., and the available Gemma Scope SAEs for Gemma 2 9B IT, at layers 10, 21, and 32 (out of 42). Our approach for detecting top latents, similar to the entity recognition method described in Section 4 focuses on the top latents with the highest minimum t-statistic score across entities, representing the most general latents. We split the dataset into train and test sets, and use the training set to select the top latents. The left panel of Figure 6 reveals a distinct separation between the latent activations at the model token when comparing correct versus incorrect responses in the test set. Using this latent as a classifier, it achieves 73.2 AUROC score, and by calibrating the decision threshold on a validation set, it gets an F1 score of 72. See Appendix P with separated errors by entity type. Table 2 illustrates the activations of the highest-scoring latent in Gemma 2B IT’s SAE on a large text corpus (Penedo et al., 2024)888https://huggingface.co/datasets/HuggingFaceFW/fineweb., showing it triggers on text related to uncertainty or undisclosed information. Figure 6 (right) illustrates the top tokens with higher logit increase by this latent, further confirming its association with concepts of unknownness.999We omit the players category since Gemma 2B IT refuses to almost all of those queries. Similar latent separations between correct and incorrect answers can also be observed in Gemma 2 9B IT (Appendix O).

### v2

@@ Abstract
Hallucinations in large language models are a widespread problem, yet the mechanisms behind whether models will hallucinate are poorly understood, limiting our ability to solve this problem. Using sparse autoencoders as an interpretability tool, we discover that a key part of these mechanisms is entity recognition, where the model detects if an entity is one it can recall facts about. Sparse autoencoders uncover meaningful directions in the representation space, these detect whether the model recognizes an entity, e.g. detecting it doesn’t know about an athlete or a movie. This suggests that models might have self-knowledge: internal representations about their own capabilities. These directions are causally relevant: capable of steering the model to refuse to answer questions about known entities, or to hallucinate attributes of unknown entities when it would otherwise refuse. We demonstrate that despite the sparse autoencoders being trained on the base model, these directions have a causal effect on the chat model’s refusal behavior, suggesting that chat finetuning has repurposed this existing mechanism. Furthermore, we provide an initial exploration into the mechanistic role of these directions in the model, finding that they disrupt the attention of downstream heads that typically move entity attributes to the final token.111We make the codebase available at https://github.com/javiferran/sae_entities.
@@ 1 Introduction
Large Language Models (LLMs) have remarkable capabilities (Radford et al., 2019; Brown et al., 2020; Hoffmann et al., 2022; Chowdhery et al., 2023) yet have a propensity to hallucinate: generating text that is fluent but factually incorrect or unsupported by available information (Ji et al., 2023; Minaee et al., 2024). This significantly limits their application in real-world settings where factuality is crucial, such as healthcare. Despite the prevalence and importance of this issue, the mechanistic understanding of whether LLMs will hallucinate on a given prompt remains limited. While there has been much work interpreting factual recall (Geva et al., 2023; Nanda et al., 2023; Chughtai et al., 2024; Yu et al., 2023), it has mainly focused on the mechanism behind recalling known facts, not on hallucinations or refusals to answer, leaving a significant gap in our understanding.
Language models can produce hallucinations due to various factors, including flawed data sources or outdated factual knowledge (Huang et al., 2023). However, an important subset of hallucinations occurs when models are prompted to generate information they don’t possess. We operationalize this phenomenon by considering queries about entities of different types (movies, cities, players, and songs). Given a question about an unknown entity, the model either hallucinates or refuses to answer. In this work, we find linear directions in the representation space that potentially encode a form of self-knowledge: assessing their own knowledge or lack thereof regarding specific entities. These directions are causally relevant for whether it refuses to answer. We note that the existence of this kind of knowledge awareness does not necessarily imply the existence of other forms of self-knowledge, and may be specific to the factual recall mechanism.
We find these directions using Sparse Autoencoders (SAEs) (Bricken et al., 2023; Huben et al., 2024). SAEs are an interpretability tool for finding a sparse, interpretable decomposition of model representations. They are motivated by the Linear Representation Hypothesis (Park et al., 2023; Mikolov et al., 2013): that interpretable properties of the input (features) such as sentiment (Tigges et al., 2023) or truthfulness (Li et al., 2023; Zou et al., 2023) are encoded as linear directions in the representation space, and that model representations are sparse linear combinations of these directions. We use Gemma Scope (Lieberum et al., 2024), which offers a suite of SAEs trained on every layer of Gemma 2 models (Team et al., 2024), and find internal representations that suggest to encode knowledge awareness in Gemma 2 2B and 9B. Additionally, we reproduce these findings for the Llama 3.1 8B model (Grattafiori et al., 2024) using LlamaScope’s SAE suite (He et al., 2024), with results presented in Appendix Q.
@@
Using sparse autoencoders (SAEs) we discover directions in the representation space on the final token of an entity, detecting whether the model can recall facts about the entity, suggesting they encode a form of knowledge awareness.
@@ 2 Sparse Autoencoders
Dictionary learning (Olshausen & Field, 1997) offers a powerful approach for disentangling features in superposition. Sparse Autoencoders (SAEs) have proven to be effective for this task (Sharkey et al., 2022; Bricken et al., 2023). SAEs project model representations into a larger dimensional space . In this work, we use the SAEs from Gemma Scope (Lieberum et al., 2024)222We use the default sparsity for each layer, the ones available in Neuronpedia (Lin & Bloom, 2024)., which use the JumpReLU SAE architecture (Rajamanoharan et al., 2024), which defines the function
@@
Equation 1 shows that the model representation can be approximately reconstructed by a linear combination of the SAE decoder latents, which often represent monosemantic features (Huben et al., 2024; Bricken et al., 2023; Templeton et al., 2024; Gao et al., 2024). By incorporating a sparsity penalty into the training loss function, we can constrain this reconstruction to be a sparse linear combination, thereby enhancing interpretability:
@@ 3 Methodology
We then categorize entities into ‘known’ or ‘unknown’. Known entities are those where the model gets at least two attributes correct, while unknown are where it gets them all wrong, we discard any in-between. To measure correctness we use fuzzy string matching.333https://github.com/seatgeek/thefuzz. See Appendix A for a description of the process. We acknowledge that this methodology might introduce some labeling inaccuracies, as the model could ‘guess’ some attributes despite not knowing about the entity or fail to recall the specific attributes we consider while knowing about the entity. However, our primary objective is to achieve a reasonable differentiation between entities rather than striving for perfect classification accuracy. Finally, we split the entities into train/validation/test (50%, 10%, 40%) sets.
@@ 4 Sparse Autoencoders Uncover Entity Recognition Directions
We also examine the level of generality of the latents by measuring their minimum separation score across entity types (): players, song, cities and movies. A high minimum separation score indicates that a latent performs robustly across entity types, suggesting strong generalization capabilities. For this purpose, for each layer () we compute , and similarly for unknown entities. The increasing trend shown in the MaxMin (red) line in Figure 2 for Gemma 2 2B and in Figure 9, Appendix D for Gemma 2 9B suggests that more generalized latents—those that distinguish between known and unknown entities across various entity types—are concentrated in these intermediate layers. This observation is replicated on Llama 3.1 8B in Appendix Q. This finding points to a hierarchical organization of entity representation within the model, with more specialized, worse quality, latents in earlier layers and more generalized, higher quality entity-type-agnostic features emerging in the middle layers.
@@
Table 1 demonstrates the activation patterns of the Gemma 2 2B topmost known entity latent on prompts with well-known entities (left), and the patterns for the topmost unknown entity latent (right), firing across entities of different types that cannot be recognized. In Appendix B we provide the activations of these latents on sentences containing a diverse set of entity types, suggesting that indeed they are highly general. To validate these latents’ reliability, we analyze their activation frequencies on 283 song titles released after the models’ knowledge cutoff (August 2024). As hypothesized, unknown entity latents show higher activation rates, while known entity latents exhibit lower activation frequencies (Appendix R). While we acknowledge potential overlap between these song titles and pre-training data, the consistent activation patterns across multiple models strengthen our confidence in these latents’ ability to distinguish between known and unknown information. In the following sections, we explore how these primary entity recognition latents influence the model’s overall behavior.
@@ 5 Entity Recognition Directions Causally Affect Knowledge Refusal
and we define a set of common knowledge refusal completions and detect if any of these occur with string matching, e.g. ‘Unfortunately, I don’t have access to real-time information…’. Gemma 2 includes both a base model, and a fine-tuned chat (i.e. instruction tuned) model. In Section 4 we found the entity recognition latents by studying the base model, but here focus on the chat model, as they have been explicitly fine-tuned to perform knowledge refusal where appropriate (Team et al., 2024)444The Gemma 2 technical report (Team et al., 2024) mentions “including subsets of data that encourage refusals to minimize hallucinations improves performance on factuality metrics”. This pattern is consistent with recent language models, such as Llama 3.1 (Dubey et al., 2024), where the explicit finetuning process for knowledge refusal has been documented., and the factuality of chat models is highly desirable.
@@
We hypothesize that entity recognition directions could be used by chat models to induce knowledge refusal. To evaluate this, we use a test set sample of 100 questions about unknown entities, and measure the number of times the model refuses by steering (as in Equation 4) with the entity recognition latents the last token of the entity and the following end-of-instruction-tokens.555We use a validation set to select an appropriate steering coefficient . In Appendix G we show generations of Gemma 2B IT with different steering coefficients. We select , which corresponds to around two times the norm of the residual stream in the layers where the entity recognition latents are present (Appendix E). Figure 3 (left) illustrates the original model refusal rate (blue bar), showing some refusal across entity types. We see that the entity recognition SAE latents found in the base model transfer to the chat model, and increasing the unknown entity latent induces almost 100% refusal across all entity types in Gemma 2 2B. Conversely, increasing the known entity latent activation slightly reduces refusal rates. We also include an Orthogonalized model baseline, which consists of doing weight orthogonalization (Arditi et al., 2024) on every matrix writing to the residual stream. Weight orthogonalization modifies each row of a weight matrix to make it perpendicular to a specified direction vector . This is achieved by subtracting the component of each row that is parallel to :
@@
By doing this operation on every output matrix in the model we ensure no component is able to write into that direction. The resulting orthogonalized model with the top unknown entity direction exhibits a large reduction in refusal responses, suggesting this direction plays a crucial role in the model’s knowledge refusal behavior. We also include the average refusal rate after steering with 10 differents random latents, using the same configuration (layer and steering coefficient) that the known and unknown entity latents respectively. Additional analysis on the Gemma 2 9B model and Llama 3.1 8B are detailed in Sections F and Q revealing similar patterns, albeit with less pronounced effects compared to the 2B model.
@@ 6 Mechanistic Analysis
In the previous section, we saw that entity recognition latents had a causal effect on knowledge refusal. Here, we look at how they affect the factual recall mechanism (aka circuit) in prompts of the format of Example 5. This has been well studied before on other language models (Nanda et al., 2023; Geva et al., 2023; Meng et al., 2022a). We replicate the approach of Nanda et al. (2023) on Gemma 2 2B and 9B and find a similar circuit. Namely, early attention heads merge the entity’s name into the last token of the entity, and downstream attention heads extract relevant attributes from the entity and move them to the final token position (Figure 4 (a, b)), this pattern holds across various entity types and model sizes (Appendix I and Appendix J). To do the analysis, we perform activation patching (Geiger et al., 2020; Vig et al., 2020; Meng et al., 2022a) on the residual streams and attention heads’ outputs (see Appendix H for a detailed explanation on the method). We use the denoising setup (Heimersheim & Nanda, 2024), where we patch representations from a clean run (with a known entity) and apply it over the run with a corrupted input (with an unknown entity).666We show the proportion of logit difference recovered after each patch in Figure 4 (a). A recovered logit difference of 1 indicates that the prediction after patching is the same as the original prediction in the clean run.
@@
Expanding on the findings of Yuksekgonul et al. (2024), who established a link between prediction accuracy and attention to the entity tokens, our study reveals a large disparity in attention between known and unknown entities, for instance the attribute extraction heads L18H5 and L20H3 (Figure 4 (c)), which are overall relevant across entity types in Gemma 2 2B (see example of attributes extracted by these heads in Appendix L). Notably, attention scores are higher when faced with a known entity. This suggests that the detected entity recognition latents might influence the attention mechanism through the ‘keys’ computation to induce this behavior. To evaluate this hypothesis we steer the residual stream with the found latents on the last token of the entity, and measure the attention scores of the entity tokens. We observe a causal relationship between the entity recognition latents and the attention patterns of the attention heads downstream, being more pronounced in the attribute extraction heads. Steering with the top unknown entity latent reduces the attention to entity, even in prompts with a known entity (Figure 4 (d)), while steering with the known entity latent increases the attention scores (Figure 4 (e)). We show in Figure 4 (f) the results of steering with a random unit vector baseline for comparison, and in Appendix K when steering with a random SAE latent. In Appendix M we illustrate the average attention score change to the entity tokens after steering on the residual streams of the last token of the entities in Gemma 2 2B, 9B, and Llama 3.1 8B with the top 3 known and unknown entity latents. The results reveal an increase/decrease attention score across upper layer heads, with the 9B model showing more subtle effects when steered using unknown latents.
@@
7 Uncertainty Directions
Our study focuses on the study of the residual streams before the answer. We hypothesize that end-of-instruction tokens, which always succeed the instruction, may aggregate information about the whole question (Marks & Tegmark, 2023).777This concept was termed by Tigges et al. (2023) as the ‘summarization motif’. We select the token model and use examples such as:
For each entity type and layer with available SAE we extract the representations of the model residual stream, for both correct and mistaken answers, and gather the SAE latent activations. We are interested in seeing whether there are SAE latents that convey information about how unsure or uncertain the model is to answer to a question, but still fails to refuse, giving rise to hallucinations. We divide the dataset of prompts into train/validation/test sets (50%, 10%, 40%).
To capture subtle variations in model uncertainty, which may be represented even when attributes are correctly recalled, we focus on quantifying differences in activation levels between correct and incorrect responses. For each latent, we compute the t-statistic in the training set using two activation samples: for correct responses and for incorrect ones. The t-statistic measures how different the two sample means are from each other, taking into account the variability within the samples:
@@ 7 Uncertainty Directions
We use a pre-trained SAE for the 13th layer (out of 18) of Gemma 2B IT888https://huggingface.co/jbloom/Gemma-2b-IT-Residual-Stream-SAEs. We note that Gemma Scope doesn’t provide SAEs for Gemma 2 2B IT., and the available Gemma Scope SAEs for Gemma 2 9B IT, at layers 10, 21, and 32 (out of 42). Our approach for detecting top latents, similar to the entity recognition method described in Section 4 focuses on the top latents with the highest minimum t-statistic score across entities, representing the most general latents. We split the dataset into train and test sets, and use the training set to select the top latents. The left panel of Figure 6 reveals a distinct separation between the latent activations at the model token when comparing correct versus incorrect responses in the test set. Using this latent as a classifier, it achieves 73.2 AUROC score, and by calibrating the decision threshold on a validation set, it gets an F1 score of 72. See Appendix P with separated errors by entity type. Table 2 illustrates the activations of the highest-scoring latent in Gemma 2B IT’s SAE on a large text corpus (Penedo et al., 2024)999https://huggingface.co/datasets/HuggingFaceFW/fineweb., showing it triggers on text related to uncertainty or undisclosed information. Figure 6 (right) illustrates the top tokens with higher logit increase by this latent, further confirming its association with concepts of unknownness.101010We omit the players category since Gemma 2B IT refuses to almost all of those queries. Similar latent separations between correct and incorrect answers can also be observed in Gemma 2 9B IT (Appendix O).